OpenStudy (anonymous):

Solve 4y - y + 11 = 38.

OpenStudy (anonymous):

4y-y=3y

OpenStudy (anonymous):

3y+11=38

OpenStudy (anonymous):

subtract 11 from both11 and 38

OpenStudy (anonymous):

Combine the y's so you get 3y+ 11=38 then you subtract 11 from 38 and both 11's cancel out so you should get 3y=27 then divide so do 27 divided by 3 and your answer is y

OpenStudy (anonymous):

y=9

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thank you never much, but why do u subtract 11 from both sides?

OpenStudy (anonymous):

So that you can get one number by itself and then the variable in this case y by itself so they can be on deprecate sides of the equation

OpenStudy (anonymous):

*seperate

OpenStudy (anonymous):

Solve 4y-y+11=38 3y+11=38 -11 -11 3y=27 /3 /3 y=9

OpenStudy (anonymous):

ok thanks for your help :-)